OpenStudy (anonymous):

medal and fan the first term of an arithmetic sequence is 2 and the fourth term is 11. find the sum of the first 50 terms.

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

3775

OpenStudy (anonymous):

i dont have any

OpenStudy (anonymous):

okay give me a sec ok

OpenStudy (anonymous):

The nth number in an arithmetic sequence = a+(n-1)d where a is the first term and d is the difference The sum of the first n numbers in an A.P. is n/2(a+nth number)

OpenStudy (anonymous):

\[\sum _{n=1}^{50} (3 n-1)=3775 \]

OpenStudy (anonymous):

Robtobey i think that's right as well that's what i got @Data_LG2

OpenStudy (anonymous):

awesome. thanks guys! everyone medal each other for a job well done! lol

OpenStudy (anonymous):

what was the answer

OpenStudy (anonymous):

the 3775

OpenStudy (anonymous):

some one medal Jessmitz

OpenStudy (anonymous):

From the Mathematica program: Table[ 3n - 1, {n, 6} ]\[\{2,5,8,11,14,17\} \]

OpenStudy (anonymous):

i was right awwwwwww yaaaa